having problems with Algebra need some help solving this problem n + 1 3/4 = 3 2/3..


Franklin

Answer by Alan3354(31532)

n + 1 3/4 = 3 2/3
n + 7/4 = 11/3
n + 21/12 = 44/12
n = 23/12